# Feedback-guided Domain Synthesis with Multi-Source Conditional Diffusion Models for Domain Generalization

### v1

@@ Abstract
Standard deep learning architectures such as convolutional neural networks and vision transformers often fail to generalize to previously unseen domains due to the implicit assumption that both source and target data are drawn from independent and identically distributed (i.i.d.) populations. In response, Domain Generalization techniques aim to enhance model robustness by simulating novel data distributions during training, typically through various augmentation or stylization strategies. However, these methods frequently suffer from limited control over the diversity of generated images and lack assurance that these images span distinct distributions. To address these challenges, we propose FDS, a novel strategy that employs diffusion models to synthesize samples from new domains by training on source distribution samples and performing domain mixing. By incorporating images that pose classification challenges to models trained on original samples, alongside the original dataset, we ensure the generation of a training set that spans a broad distribution spectrum. Our comprehensive evaluations demonstrate that this methodology sets new benchmarks in domain generalization performance across a range of challenging datasets, effectively managing diverse types of domain shifts. The implementation is available at: https://github.com/Mehrdad-Noori/FDS.git.
@@ 1 Introduction
Domain generalization (DG) [9] aims to solve the issue of domain shift by training models using data from one or more source domains, so they perform well out-of-the-box on new, unseen domains. Recently, various techniques have been developed to tackle this problem [10, 11], including domain aligning [12, 13, 14], meta-learning [15, 16], data augmentation [17, 18, 19], ensemble learning[20], applying self-supervised learning techniques [21, 22], and adding regularization methods [23, 24]. These strategies are designed to make models more adaptable and capable of handling data that they were not explicitly trained on, making them more useful in real-world situations where the exact nature of future data cannot be predicted. Among these techniques, a notable category focuses on synthesizing samples from different distributions to mimic target distributions. This is achieved through strategies like image transformation [17, 25], style transfer [26, 27], learnable augmentation networks [28, 29, 30], and feature-level stylization [20, 31]. However, many of these methods face challenges in controlling the synthesis process, often resulting in limited diversity where primarily only textures are altered.
@@
Figure 1: Bridging Domain Gaps with FDS: Comprehensive Distribution Coverage Across Domains.
@@
In this study, we introduce an innovative approach using diffusion model to address the challenge of domain generalization. Known for their exceptional ability to grasp intricate distributions and semantics, diffusion models excel at producing high-quality, realistic samples [32, 33, 34, 35]. We exploit this strength by training a diffusion model that is conditioned on various classes and domains present in the training dataset, aiming to master the distribution of source domains. Through the process of domain mixing during generation, we create images that appear to originate from new, pseudo-target domains. To ensure the development of a robust classifier, we initially select generated samples that are difficult for a model trained solely on the original source domains to classify. Subsequently, we train the model using a combination of these challenging images and the original dataset. In this manner, we ensure that our model is fed with images of the widest possible diversity, thereby significantly enhancing its ability to generalize across unseen domains. Our contributions can be summarized as follows:
@@
  • We introduce a novel DG approach that leverages diffusion models conditioned on different domains and classes to generate samples from novel, pseudo-target domains through domain interpolation. This approach increases the diversity and realism of the generated images;
@@
  • We propose an innovative strategy for selecting images that pose a challenge to a classifier trained on original images, ensuring the diversity of the final sample set. We then train a classifier on this enriched dataset, comprising both challenging generated images and original images, to enhance its generalization capabilities;
@@
  • We conduct extensive experiments across various benchmarks to validate the effectiveness of our method. These experiments demonstrate FDS’s ability to significantly improve the robustness and generalization of models across a wide range of unseen domains, achieving state-of-art performance on these datasets.
The rest of this paper is structured as follows. The next section reviews previous work on domain generalization and recent uses of diffusion models. Section 4 defines the DG problem we focus on and introduces our proposed method. Section 5 presents the experimental setup employed to validate our method. In Section 6, we share the results and findings of our experiments, showing how our method improves upon previous work. We wrap up the paper in Section 7, discussing our main results and ideas for what could be studied next.
@@ 2.1 Domain Generalization (DG)
Domain Generalization, a concept first introduced by Blanchard et al. in 2011 [9], has seen growing interest in the field of computer vision. This interest has spurred the development of a broad spectrum of methods aimed at enabling models to generalize across unseen domains. These methods can be categorized into several key approaches, each targeting the DG problem from a unique angle.
Domain alignment.  Strategies based on domain alignment focus on minimizing the disparity between source domains by learning representations that are invariant across these domains. Various techniques, such as moment matching [36], discriminant analysis [12], minimizing maximum mean discrepancy (MMD)[37], employing contrastive loss[38], and domain-adversarial learning [39], are employed to achieve this goal.
Meta-learning.  Approaches using meta-learning address DG [15, 16] by considering a bi-level optimization problem, where the model is fine-tuned on meta-source domains to minimize the error on a meta-target domain. This is often achieved through episodic training that may involve updating all [15] or a subset [16] of the network’s parameters.
Data augmentation.  A key strategy in DG, data augmentation aims to enhance model robustness against domain shifts encountered during deployment. This objective is pursued through a variety of techniques, such as learnable augmentation, off-the-shelf style transfer, and augmentation at the feature level. Learnable augmentation models utilizes networks to create images from training data, ensuring their distribution diverges from that of the source domains [28, 30, 29]. Meanwhile, off-the-shelf style transfer based methods seek to transform the appearance of images from one domain to another or modify their stylistic elements, often through Adaptive Instance Normalization (AdaIN)[40, 26]. Unlike most augmentation approaches that modify pixel values, some propose altering features directly, a technique inspired by the finding that CNN features encapsulate style information[41, 20] in their statistics.
Ensemble learning.  Techniques based on this approach focus on improving robustness to out-of-distribution (OOD) data by developing multiple models tailored to specific domains [20]. The collective predictions of these models, weighted by either the resemblance of a target sample to source domains or the confidence levels of the models, offer a nuanced assessment for target domain samples [42].
Self-supervised learning.  Conversely, self-supervised learning methods concentrate on fostering domain-agnostic representations through the pre-training of models on unsupervised tasks, such as puzzle solving [21, 43], image reconstruction with autoencoders [44], or engaging in contrastive learning [45].
Disentangled representation learning  Methods in this category try to segregate features into those unique to specific domains and those common across all domains [46, 47]. This segregation is facilitated by domain-specific masks that highlight the features most pertinent to images from the target domain [47].
Regularization  Lastly, regularization methods in DG are designed to refine the training process, encouraging the development of features that are universally applicable across domains [24, 48, 49, 50]. These methods typically build on the empirical risk minimization (ERM) framework [51], incorporating additional objectives such as distillation [48, 50], stochastic weight averaging [24], or distributionally robust optimization (DRO) [49] to promote generalization.
@@ 2.2 Diffusion Models
Diffusion models have recently surpassed Generative Adversarial Networks (GANs) as the leading technique for image synthesis. Innovations like denoising diffusion probabilistic models (DDPMs) [33] and denoising diffusion implicit models (DDIMs) [35] have significantly sped up the image generation process. Rombach et al. [34] introduced latent diffusion models (LDMs), also known as stable diffusion models, which enhance both training and inference efficiency while facilitating text-to-image and image-to-image conversions. Following this advancement, various enhancements and extensions have been developed, including stable diffusion XL (SDXL) [52] and ControlNet [53]. These models introduce the capability to guide the generation process using additional inputs, such as depth or semantic information, allowing for more controlled and versatile image creation.
Despite the impressive capabilities of diffusion models in generating images, their application in domain generalization has been minimally explored. Building on the concept of using style transfer networks, Yu et al. have applied diffusion models to adapt the style of target images onto source images [54], though this approach incurs additional computational costs during deployment. In a recent development, Niemeijer et al. introduced DIDEX [55], which leverages textual inputs to alter the style of input images, thereby generating a pseudo-target dataset that enhances model performance through adaptation to these pseudo-target samples within the unsupervised domain adaptation (UDA) framework. However, this method is constrained by the need for human intervention in selecting text inputs that align with the anticipated characteristics of the target domain. Diverging from the strategy of stylizing input images, our method adopts a unique approach by directing the diffusion process to produce new images from novel, diverse distributions while preserving the semantic integrity of the source samples.
@@ 3 Theoretical Motivation
Our classifier, represented by $f$ with parameters $\theta$, aims to craft a unified model from $n$ source domains $\{\mathcal{D}^{1},\ldots,\mathcal{D}^{n}\}$ that adapts to a novel target domain $\mathcal{D}^{\mathcal{T}}$. Within any domain $\mathcal{D}$, we measure classification loss by:
@@
where $x$ and $y$ denote the input and its corresponding label, respectively, and $\ell(f(x;\theta),y)$ is the cross entropy loss.
@@
Here, $|\mathcal{D}^{i}|$ counts the data points in domain $i$, with $x_{i}^{k}$ as the $k$-th point and $y_{i}^{k}$ as its label. However, the accuracy of ERM-trained models drops when data shifts occur across domains due to inadequate OOD generalization. To enhance ERM, Chapelle et al. introduced Vicinal Risk Minimization (VRM) [57], which substitutes point-wise estimates with density estimation in the vicinity distribution around each observation within each domain. Practical implementation often involves data augmentation to introduce synthetic samples from these density estimates. Traditional augmentation processes a single data point $\tilde{x}_{i}^{k}$ from domain $i$ to yield $\tilde{x}_{i}^{k}=g(\tilde{x}_{i}^{k})$, where $g(\cdot)$ denotes a basic transformation and $\tilde{x}_{i}^{k}$ is the modified data point.
@@ 4 Method
Figure 2: Overview of the proposed architecture for FDS. (left) Multi-Source Training of Diffusion Model Conditioned on Class and Domain of the training images. (right) generating pseudo-novel image set using proposed interpolation and filtering mechanism of FDS.
To advance the generalization ability of our classifier, we aim to utilize generated image samples that traverse domain gaps yet retain class semantics. This objective is realized through a three-step methodology:
  • Step 1: We begin by training an image generator that masters the class-specific distributions within each domain, enabling the synthesis of class-consistent samples within those domains.
  • Step 2: Then, we introduce a strategy for generating inter-domain images to span the domain gaps effectively.
  • Step 3: Last, we filter overly simplistic samples from synthetic inter-domain images and train the final model on this refined set alongside original domain images for enhanced OOD generalization.
In the follow subsections, we present each of these steps in details.
@@ 4.1 Image Generator
This approach enables a single diffusion model to understand and generate images across different domains for each class, using text-based conditions (prompts). By dynamically adjusting conditions, the model efficiently learns varied representations without needing multiple models for each domain. Thus, we employ a textual template $c(x)$, denoting “[y], [D]” where y is the class label and D is the domain name of the input $x$, and proceed to train our diffusion model using this template across all images from the source domains. Upon completing training, we possess the capability to create a new sample for class $y^{k}$, belonging to the set $\{y^{1},\ldots,y^{m}\}$, within domain $\mathcal{D}^{i}$. This is achieved by decoding a denoised representation after $t$ timesteps, $\tilde{x}_{t}^{i,k}=\tilde{\mathcal{E}}(\Phi_{t}(\mathcal{D}^{i},y^{k}))$, where $\Phi_{t}(\mathcal{D}^{i},y^{k})$ is the denoised representation that originates from random Gaussian noise conditioned on $\mathcal{D}^{i}$ and $y^{k}$.
@@ 4.2 Domain Mixing
As illustrated in 1, In order to leverage image generator to synthesize images from new unseen distributions we propose two mixing strategies:
4.2.1 Noise Level Interpolation.
@@
4.2.2 Condition Level Interpolation.
@@ 5 Experimental Setup
Table 1: Comparison to the state-of-art on three benchmarks, reporting the mean and standard deviation across three runs.The best results are in bold fonts.
@@
Method Category PACS VLCS OfficeHome Avg. ERM [51] Baseline 85.5 $\pm$0.2 77.5 $\pm$0.4 66.5 $\pm$0.3 76.50 $\pm$0.3 Mixstyle [20] Data Augmentation 85.2 $\pm$0.3 77.9 $\pm$0.5 60.4 $\pm$0.3 74.50 $\pm$0.4 RSC [23] 85.2 $\pm$0.9 77.1 $\pm$0.5 65.5 $\pm$0.9 75.93 $\pm$0.8 Mixup [60] 84.6 $\pm$0.6 77.4 $\pm$0.6 68.1 $\pm$0.3 76.70 $\pm$0.5 SagNet [61] 86.3 $\pm$0.2 77.8 $\pm$0.5 68.1 $\pm$0.1 77.40 $\pm$0.3 DCAug [59] 86.1 $\pm$0.7 78.6 $\pm$0.4 68.3 $\pm$0.4 77.66 $\pm$0.5 IRM [62] Domain-Invariant 83.5 $\pm$0.8 78.5 $\pm$0.5 64.3 $\pm$2.2 75.43 $\pm$0.5 GroupDRO [49] 84.4 $\pm$0.8 76.7 $\pm$0.6 66.0 $\pm$0.7 75.70 $\pm$0.7 DANN [63] 83.6 $\pm$0.4 78.6 $\pm$0.4 65.9 $\pm$0.6 76.03 $\pm$0.5 MMD [64] 84.6 $\pm$0.5 77.5 $\pm$0.9 66.3 $\pm$0.1 76.13 $\pm$0.5 SelfReg [45] 86.5 $\pm$0.3 77.5 $\pm$0.0 69.4 $\pm$0.2 77.80 $\pm$0.2 CORAL [65] 86.2 $\pm$0.3 78.8 $\pm$0.6 68.7 $\pm$0.3 77.90 $\pm$0.4 mDSDI [66] 86.2 $\pm$0.2 79.0 $\pm$0.3 69.2 $\pm$0.4 78.13 $\pm$0.3 ERM(reproduced[51] 84.30 $\pm$1.1 76.23 $\pm$1.1 64.60 $\pm$1.1 75.04 $\pm$0.8 ERM + FDS (ours) 8 7.38 $\pm$1.6 79.46 $\pm$1.6 69.17 $\pm$1.1 78.67 $\pm$1.4 SWAD [24] Ensemble Learning 88.1 $\pm$0.1 79.1 $\pm$0.1 70.6 $\pm$0.2 79.27 $\pm$0.1 TeachDCAug [59] 88.4 $\pm$0.2 78.8 $\pm$0.4 70.4 $\pm$0.2 79.20 $\pm$0.3 DNA [67] 88.4 $\pm$0.1 79.0 $\pm$0.1 71.2 $\pm$0.1 79.53 $\pm$0.1 DIWA [68] 88.8 $\pm$0.4 79.1 $\pm$0.2 71.0 $\pm$0.1 79.63 $\pm$0.2 SWAD (reproduced[24] 88.11 $\pm$0.4 78.87 $\pm$0.5 70.34 $\pm$0.4 79.11 $\pm$0.4 SWAD + FDS (ours) 90.50 $\pm$0.3 79.73 $\pm$0.5 73.51 $\pm$0.4 81.25 $\pm$0.4
5.1 Dataset
Following Gulrajani and Lopez-Paz [51], we compare our proposed approach to the current state-of-art using three challenging datasets with different characteristics: PACS [46], VLCS [69], and OfficeHome [70].
@@ 5.1 Dataset
The PACS dataset [46] has a total of 9,991 photos divided into four distinct domains, $d\in$Art, Cartoon, Photo, Sketch} and seven distinct classes. The second dataset, VLCS, comprises 10,729 photos from four separate domains, $d\in$Caltech101, LabelMe , SUN09, VOC2007} and five different classes. The third dataset, OfficeHome, includes a total of 15,588 photos taken from four domains, $d\in$Art, Clipart, Product, Real}, and 65 classes.
5.2 Implementation Details
To ensure a fair comparison, we adopt the DomainBed framework [51], a comprehensive benchmark that encompasses prominent domain generalization (DG) methodologies under a uniform evaluation protocol. Following this framework, we employ a leave-one-out strategy for DG dataset assessment where one domain serves as the test set while the others form the training set. A subset of the training data, constituting 20%, is designated as the validation set [^1] The model demonstrating peak accuracy on this validation set is selected as optimal. This model undergoes evaluation on the test domain, providing the unseen domain accuracy.. The aggregate result for each dataset represents the mean accuracy derived from varying the test domain. To ensure reliability, experiments are replicated three times, each with a unique seed.
Moreover, to rigorously test our method, we base our classifiers on the SWAD technique [51], a stable ensemble approach surpassing several recent DG methods. Our method’s efficacy is showcased without extensive hyperparameter tuning, adhering to SWAD’s recommended settings.
@@ 5.2 Implementation Details
For image synthesis, the original Stable Diffusion framework is utilized with $\mathrm{DDIM}\!=\!50$ steps. The PACS and VLCS datasets prompt the generation of $N\!=\!32,000$ samples per class, whereas OfficeHome, with its 65 classes, necessitates $N\!=\!16,000$ samples. Image generation spans an interpolation range of $\alpha\!\in\![0.3,0.7]$ and a Noise Level Interpolation range of $\mathcal{T}\!\in\![20,45]$. This diversified parameter selection, rather than optimizing hyperparameters per dataset, acknowledges each domain’s unique shift. Our filtering mechanism then identifies the most informative images from the generated pool. The selection of an optimal dataset size, $N_{L}$, underwent extensive exploration, detailed further in Section 6.2.

### v2

# FDS: Feedback-guided Domain Synthesis with Multi-Source Conditional Diffusion Models for Domain Generalization

@@ Abstract
Domain Generalization techniques aim to enhance model robustness by simulating novel data distributions during training, typically through various augmentation or stylization strategies. However, these methods frequently suffer from limited control over the diversity of generated images and lack assurance that these images span distinct distributions. To address these challenges, we propose FDS, Feedback-guided Domain Synthesis, a novel strategy that employs diffusion models to synthesize novel, pseudo-domains by training a single model on all source domains and performing domain mixing based on learned features. By incorporating images that pose classification challenges to models trained on original samples, alongside the original dataset, we ensure the generation of a training set that spans a broad distribution spectrum. Our comprehensive evaluations demonstrate that this methodology sets new benchmarks in domain generalization performance across a range of challenging datasets, effectively managing diverse types of domain shifts. The implementation is available at: https://github.com/Mehrdad-Noori/FDS.git.
@@ 1 Introduction
Domain generalization (DG) [9] aims to solve the issue of domain shift by training models using data from one or more source domains, so they perform well out-of-the-box on new, unseen domains. Recently, various techniques have been developed to tackle this problem [10, 11], including domain aligning [12, 13, 14], meta-learning [15, 16], data augmentation [17, 18, 19], ensemble learning[20], self-supervised learning [21, 22], and regularization methods [23, 24]. These strategies are designed to make models more adaptable and capable of handling data that they were not explicitly trained on, making them more useful in real-world situations where the exact nature of future data cannot be predicted. Among these techniques, a notable category focuses on synthesizing samples from different distributions to mimic target distributions. This is achieved through strategies like image transformation [17, 25], style transfer [26, 27], learnable augmentation networks [28, 29, 30], and feature-level stylization [20, 31]. However, many of these methods face challenges in controlling the synthesis process, often resulting in limited diversity where primarily only textures are altered.
@@
Figure 1: Generating new, pseudo-domains with FDS: Comprehensive distribution coverage from domain $D_{1}$ to $D_{2}$.
@@
In this study, we introduce an innovative approach using diffusion model, named Feedback-guided Domain Synthesis (FDS), to address the challenge of domain generalization. Known for their exceptional ability to grasp intricate distributions and semantics, diffusion models excel at producing high-quality, realistic samples [32, 33, 34, 35]. We exploit this strength by training a single diffusion model that is conditioned on various domains and classes present in the training dataset, aiming to master the distribution of source domains. As illustrated in Figure 1, through the process of domain interpolation and mixing during generation, we create images that appear to originate from novel, pseudo-domains. To ensure the development of a robust classifier, we initially select generated samples that are difficult for a model trained solely on the original source domains to classify. Subsequently, we train the model using a combination of these challenging images and the original dataset. In this manner, we ensure that our model is fed with images of the widest possible diversity, thereby significantly enhancing its ability to generalize across unseen domains. Our contributions can be summarized as follows:
@@
  • We introduce a novel DG approach that leverages diffusion models conditioned on multiple domains and classes to generate samples from novel, pseudo-domains through domain interpolation. This approach increases the diversity and realism of the generated images;
@@
  • We propose an innovative strategy for selecting images that pose a challenge to a classifier trained on original images, ensuring the diversity of the final sample set. By incorporating these challenging images with the original dataset, we ensure a comprehensive and diverse training set, significantly improving the model’s generalization capabilities;
@@
  • We conduct extensive experiments across various benchmarks and perform different analyses to validate the effectiveness of our method. These experiments demonstrate FDS’s ability to significantly improve the robustness and generalization of models across a wide range of unseen domains, achieving SOTA performance.
@@ 2.1 Domain Generalization (DG)
Domain Generalization, a concept first introduced by Blanchard et al. in 2011 [9], has seen growing interest in the field of computer vision. This interest has spurred the development of a broad spectrum of methods aimed at enabling models to generalize across unseen domains. These include approaches based on domain alignment like moment matching [36], discriminant analysis [12] and domain-adversarial learning [37], meta-learning approaches [15, 16] which solve a bi-level optimization problem where the model is fine-tuned on meta-source domains to minimize the error on a meta-target domain, ensemble learning techniques [20, 38] improving robustness to OOD data by training multiple models tailored to specific domains, self-supervised learning methods fostering domain-agnostic representations through the pre-training of models on unsupervised tasks [21, 39, 40] or via contrastive learning [41], approaches leveraging disentangled representation learning [42, 43] segregating domain-specific features from those common across all domains, and regularization methods which build on the empirical risk minimization (ERM) framework [44], incorporating additional objectives such as distillation [45, 46], stochastic weight averaging [24], or distributionally robust optimization (DRO) [47] to promote generalization.
A key strategy in DG, data augmentation aims to enhance model robustness against domain shifts encountered during deployment. This objective is pursued through a variety of techniques, such as learnable augmentation, off-the-shelf style transfer, and augmentation at the feature level. Learnable augmentation models utilizes networks to create images from training data, ensuring their distribution diverges from that of the source domains [28, 30, 29]. Meanwhile, off-the-shelf style transfer based methods seek to transform the appearance of images from one domain to another or modify their stylistic elements, often through Adaptive Instance Normalization (AdaIN)[48, 26]. Unlike most augmentation approaches that modify pixel values, some propose altering features directly, a technique inspired by the finding that CNN features encapsulate style information[49, 20] in their statistics.
@@ 2.2 Diffusion Models
Diffusion models have recently surpassed Generative Adversarial Networks (GANs) as the leading technique for image synthesis. Innovations like denoising diffusion probabilistic models (DDPMs) [33] and denoising diffusion implicit models (DDIMs) [35] have significantly sped up the image generation process. Rombach et al. [34] introduced latent diffusion models (LDMs), also known as stable diffusion models, which enhance both training and inference efficiency while facilitating text-to-image and image-to-image conversions. Extensions of these models, such as stable diffusion XL (SDXL) [50] and ControlNet [51], have been developed to further guide the generation process with additional inputs, such as depth or semantic information, allowing for more controlled and versatile image creation. Recent studies have shown that augmenting datasets using diffusion models can improve performance in general vision tasks [52, 53].
Despite the capabilities of diffusion models in generating images, their application in domain generalization has been minimally explored. Yue et al. [54] proposed a method called DSI to address OOD prediction by transforming testing samples back to the training distribution using multiple diffusion models, each trained on a single source distribution. While effective, DSI requires multiple models during prediction, making it impractical for real-time deployment. In contrast, our method only uses diffusion during training to synthesize pseudo-domains and select more challenging images, keeping the computational complexity the same during testing while achieving significantly better performance.
In another study, CDGA [55] uses diffusion models to generate synthetic images that fill the gap between different domain pairs through a simple interpolation. While effective, CDGA relies on generating a large number of synthetic images (e.g., 5M for PACS) and employs naive interpolation. Additionally, CDGA generates samples with additional text descriptions that are not provided in standard DG benchmarks. In contrast, our method employs novel and more efficient interpolation techniques and a unique filtering mechanism that selects challenging images. We demonstrate that both of these innovations are crucial and significantly effective for improving generalization.
Figure 2: Overview of the proposed architecture for FDS. (top) Multi-source training of diffusion model conditioned on class and domain of the training images. (bottom) Generating novel pseudo-domain using the proposed interpolation and filtering mechanism of FDS.
@@ 3 Theoretical Motivation
Our classifier, represented by $f$ with parameters $\theta$, aims to craft a unified model from $n$ source domains $\{\mathcal{D}^{1},\ldots,\mathcal{D}^{n}\}$ that adapts to a novel target domain $\mathcal{D}^{\mathcal{T}}$. Within any domain $\mathcal{D}$, we measure classification loss by
@@
where $x$ and $y$ denote the input and its corresponding label, respectively, and $\ell(f(x;\theta),y)$ is the cross entropy loss in this work.
@@
Here, $|\mathcal{D}^{i}|$ counts the number of samples in domain $i$, with $x_{i}^{k}$ as the $k$-th sample and $y_{i}^{k}$ its label. However, the accuracy of ERM-trained models drops when data shifts occur across domains due to inadequate OOD generalization. To enhance ERM, Chapelle et al. introduced Vicinal Risk Minimization (VRM) [57], which substitutes point-wise estimates with density estimation in the vicinity distribution around each observation within each domain. Practical implementation often involves data augmentation to introduce synthetic samples from these density estimates. Traditional augmentation processes a single data point $\tilde{x}_{i}^{k}$ from domain $i$ to yield $\tilde{x}_{i}^{k}=g(\tilde{x}_{i}^{k})$, where $g(\cdot)$ denotes a basic transformation and $\tilde{x}_{i}^{k}$ is the modified data point.
@@ 4 Method
To advance the generalization ability of our classifier, we aim to utilize generated image samples that traverse domain gaps yet retain class semantics. This objective is realized through a three-step methodology of our proposed method FDS. Step 1: we begin by training an image generator that masters the class-specific distributions of all source domains, enabling the synthesis of class-consistent samples within those domains. Step 2: we then introduce a strategy for generating inter-domain images to span the domain gaps effectively. Step 3: last, we filter overly simplistic samples from synthetic inter-domain images and train the final model on this refined set alongside original domain images for enhanced OOD generalization. The process is illustrated in Figure 2.
@@ 4.1 Image Generator
This approach enables a single diffusion model to understand and generate images across different domains for each class, using text-based conditions (prompts). By dynamically adjusting conditions, the model efficiently learns varied representations without needing multiple models for each domain. Thus, we employ a textual template $c(x)$, denoting “[y], [D]” where y is the class label and D is the domain name of the input $x$, and proceed to train our diffusion model using this template across all images from the source domains. Upon completing training, we can create a new sample for class $y^{k}$, belonging to the set $\{y^{1},\ldots,y^{m}\}$, within domain $\mathcal{D}^{i}$. This is achieved by decoding a denoised representation after $t$ timesteps, $\tilde{x}_{t}^{i,k}=\tilde{\mathcal{E}}(\Phi_{t}(\mathcal{D}^{i},y^{k}))$, where $\Phi_{t}(\mathcal{D}^{i},y^{k})$ is the denoised representation that originates from random Gaussian noise conditioned on $\mathcal{D}^{i}$ and $y^{k}$.
@@ 4.2 Domain Mixing
We propose two mixing strategies to synthesize images from new, pseudo distributions, based on noise-level interpolation and condition-level interpolation.
4.2.1 Noise Level Interpolation
@@
4.2.2 Condition Level Interpolation
@@ 5 Experimental Setup
Table 1: Leave-one-out accuracy (%) results on the PACS, VLCS, and OfficeHome benchmarks. "Aug." indicates whether advanced augmentation or domain mixing techniques are used. The best results and second-best results are highlighted.
@@
Method Aug. PACS VLCS Office  Avg. Standard Methods ERM (baseline[44] 85.5 $\pm$0.2 77.5 $\pm$0.4 66.5 $\pm$0.3 76.5 ERM (reproduced) 84.3 $\pm$1.1 76.2 $\pm$1.1 64.6 $\pm$1.1 75.0 IRM [60] 83.5 $\pm$0.8 78.5 $\pm$0.5 64.3 $\pm$2.2 75.4 GroupDRO [47] 84.4 $\pm$0.8 76.7 $\pm$0.6 66.0 $\pm$0.7 75.7 Mixup [61] 84.6 $\pm$0.6 77.4 $\pm$0.6 68.1 $\pm$0.3 76.7 CORAL [62] 86.2 $\pm$0.3 78.8 $\pm$0.6 68.7 $\pm$0.3 77.9 MMD [63] 84.6 $\pm$0.5 77.5 $\pm$0.9 66.3 $\pm$0.1 76.1 DANN [64] 83.6 $\pm$0.4 78.6 $\pm$0.4 65.9 $\pm$0.6 76.0 SagNet [65] 86.3 $\pm$0.2 77.8 $\pm$0.5 68.1 $\pm$0.1 77.4 RSC [23] 85.2 $\pm$0.9 77.1 $\pm$0.5 65.5 $\pm$0.9 75.9 Mixstyle [20] 85.2 $\pm$0.3 77.9 $\pm$0.5 60.4 $\pm$0.3 74.5 mDSDI [66] 86.2 $\pm$0.2 79.0 $\pm$0.3 69.2 $\pm$0.4 78.1 SelfReg [41] 85.6 $\pm$0.4 77.8 $\pm$0.9 67.9 $\pm$0.7 77.1 DCAug [59] 86.1 $\pm$0.7 78.6 $\pm$0.4 68.3 $\pm$0.4 77.7 CDGA [55] 88.5 $\pm$0.5 79.6 $\pm$0.3 68.2 $\pm$0.6 78.8 ERM + FDS (ours) 88.8 $\pm$0.1 79.8 $\pm$0.5 71.1 $\pm$0.1 79.9 WA Methods SWAD (baseline[24] 88.1 $\pm$0.1 79.1 $\pm$0.1 70.6 $\pm$0.2 79.3 SWAD (reproduced) 88.1 $\pm$0.4 78.9 $\pm$0.5 70.3 $\pm$0.4 79.1 SelfReg SWA [41] 86.5 $\pm$0.3 77.5 $\pm$0.0 69.4 $\pm$0.2 77.8 DNA [67] 88.4 $\pm$0.1 79.0 $\pm$0.1 71.2 $\pm$0.1 79.5 DIWA [68] 88.8 $\pm$0.4 79.1 $\pm$0.2 71.0 $\pm$0.1 79.6 TeachDCAug [59] 88.4 $\pm$0.2 78.8 $\pm$0.4 70.4 $\pm$0.2 79.2 SWAD + FDS (ours) 90.5 $\pm$0.3 79.7 $\pm$0.5 73.5 $\pm$0.4 81.3
@@
Datasets.  Following [44], we compare our proposed approach to the current state-of-art using three challenging datasets - 12 individual target domains - with different characteristics: PACS [42], VLCS [69], and OfficeHome [70]. The PACS dataset has a total of 9,991 photos divided into four distinct domains, $d\in$Art, Cartoon, Photo, Sketch} and seven distinct classes. The second dataset, VLCS, comprises 10,729 photos from four separate domains, $d\in$Caltech101, LabelMe , SUN09, VOC2007} and five different classes. The third dataset, OfficeHome, includes a total of 15,588 photos taken from four domains, $d\in$Art, Clipart, Product, Real}, and 65 classes.
@@
Implementation Details.  To ensure a fair comparison, we adopt the DomainBed framework [44], a comprehensive benchmark that encompasses prominent domain generalization (DG) methodologies under a uniform evaluation protocol. Following this framework, we employ a leave-one-out strategy for DG dataset assessment where one domain serves as the test set while the others form the training set. A subset of the training data, constituting 20%, is designated as the validation set [^1] The model with peak accuracy on this validation set is selected for evaluation on the test domain, providing unseen domain accuracy.. The aggregate result for each dataset represents the mean accuracy derived from varying the test domain. To ensure reliability, experiments are replicated three times, each with a unique seed. Moreover, to rigorously test our method, we try both ERM and SWAD classifiers with FDS. The former is considered a baseline in standard training methods, while the latter serves as a baseline for weight averaging (WA) methods. SWAD is essentially ERM but with weight averaging applied during training using multiple steps based on the validation set. To have a fair comparison with other methods, for the ERM baseline, we strictly follow the hyperparameter tuning proposed in [24]. For SWAD, as in the original work, we did not tune any parameters and used the default values. For image synthesis, the original Stable Diffusion framework [34] is utilized with $\mathrm{DDIM}\!=\!50$ steps. The PACS and VLCS datasets prompt the generation of $N\!=\!32,000$ samples per class, whereas OfficeHome, with its 65 classes, necessitates $N\!=\!16,000$ samples. Image generation spans an interpolation range of $\alpha\!\in\![0.3,0.7]$ and a Noise Level Interpolation range of $\mathcal{T}\!\in\![20,45]$. This diversified parameter selection, rather than optimizing hyperparameters per dataset, acknowledges each domain’s unique shift. Our filtering mechanism then identifies the most informative images from the generated pool. The selection of an optimal dataset size, $N_{L}$, underwent extensive exploration, detailed further in Section 6.2.